Six pears and three apples cost $3.90. Two pears and five apples cost $3.30. How much does one pear cost?
Answer by checkley75(3666)

6P+3A=3.90
2P+5A=3.30 MULTIPLY THIS EQUATION BY -3 & ADD
-6P-15A=-9.90
6P+3A=3.90
----------------------
-12A=-6.00
A=-6.00/-12
A=.50 PRICE OF 1 APPLE.
6P+3*.50=3.90
6P+1.50=3.90
6P=3.90-1.50
6P=2.40
P=2.40/6
P=.40 PRICE OF 1 PEAR.
PROOF
2*.40+5*.50=3.30
.80+2.50=3.30
3.30=3.30